# MetaTrading: An Immersion-Aware Model Trading Framework for Vehicular Metaverse Services

### v2

@@ IV-C Multi-MSP Multi-MU Game as EPEC
As shown in Fig. 6, MUs and MSPs negotiate allocation strategies of computational and communication resources and reward decisions to maximize their benefits. In the upper level, MSPs determine the rewards they are willing to offer by considering their cost, the responses of MUs, and the decisions of other MSPs. In the lower level, each MU $m$ gives the optimal allocation response for computational and communication resources by considering its resource constraints, cost, and rewards from different MSPs. The objective of EPEC is to find the equilibria at two levels, i.e., the point at which the MSPs’ (leaders’) utilities are maximized given that the MUs (followers) will choose their best responses. For the proposed EPEC, the equilibria at two levels are defined as follows.
@@
$\Psi_{n}\left(\bm{f}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{B}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{p}_{n}^{*},\bm{p}_{-n}^{*}\right)\geq$
$\Psi_{n}\left(\bm{f}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{B}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{p}_{n},\bm{p}_{-n}^{*}\right),\forall n\in\mathcal{N},$
where $\bm{p}_{-n}^{*}$ denotes the optimal rewards offered by other MSPs.
@@
$\max_{p_{n}}\,\,\Psi_{n}=\psi_{n}-\sum_{m=1}^{M}\bigg{[}p_{mn}I_{mn}\bigg{(}\frac{1}{2}\tau_{n}$ (14)
$-\frac{x_{m}\tau_{n}\log(1/\theta_{m})}{f_{mn}^{*}}-\frac{b_{mn}}{B_{mn}^{*}\log_{2}(1+\varsigma_{mn})}\bigg{)}\bigg{]},$
[…]
@@ V-A Lower Level: Optimal Resource Allocation for MUs
Definition 3.
A maximization problem has a unique global optimum if the following conditions are satisfied [37]: (1) The objective function is strictly concave; (2) All constraints (equality and inequality) are affine; and (3) The strategy space is compact and convex.
@@
We find that $\bm{H}^{f}_{m}$ and $\bm{H}^{B}_{m}$ are both negative semi-definite. Hence, $\bm{H}_{m}$ is negative semi-definite, implying $\Phi_{m}(\bm{f}_{m},\bm{B}_{m})$ is strictly concave. Since the constraints of Problem 1 are affine and the strategy space is closed and bounded, Definition 3 ensures the existence of a unique global optimum $(\bm{f}_{m}^{*},\bm{B}_{m}^{*})$. ∎
Consequently, MU $m$ has a unique best-response resource allocation for each MSP by solving Problem 1. From the Karush–Kuhn–Tucker (KKT) conditions, we establish the following proposition.
@@
$f_{mn}^{*}=\begin{cases}\sqrt{\frac{p_{mn}I_{mn}x_{m}\tau_{n}}{c_{m}^{f}}},&\text{ $p_{mn}\geq\frac{F^{2}}{I_{mn}}$},\\ F\sqrt{\frac{x_{m}\tau_{n}}{c_{m}^{f}}},&\text{\,\,$0<p_{mn}<\frac{F^{2}}{I_{mn}}$},\\ 0,&\,\,\text{otherwise},\end{cases}$ (18)
@@
$B_{mn}^{*}=\begin{cases}\sqrt{\frac{p_{mn}I_{mn}b_{mn}}{c_{m}^{B}\log_{2}(1+\varsigma_{mn})}},&\text{$p_{mn}\geq\frac{F^{2}}{I_{mn}}$},\\ F\sqrt{\frac{b_{mn}}{c_{m}^{B}\log_{2}(1+\varsigma_{mn})}},&\text{\,\,$0<p_{mn}<\frac{F^{2}}{I_{mn}}$},\\ 0,&\text{otherwise},\end{cases}$ (19)
where $F=\log(1/\theta_{m})\sqrt{\frac{x_{m}c_{m}^{f}}{\tau_{n}}}+\frac{\sqrt{b_{mn}c_{m}^{B}}}{\tau_{n}\sqrt{\log_{2}(1+\varsigma_{mn})}}$. For the detailed proof of Proposition 1, please refer to Appendix A.
@@ V-B Upper Level: Optimal Reward Equilibrium among MSPs
There exists a unique Nash equilibrium in the multi-MSP rewarding game $\Omega$, ensuring a unique optimal set of reward decisions $\{\bm{p}_{n}^{*}\}_{n\in\mathcal{N}}$.
@@
The proof of $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}<0$ and $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}\partial p_{m^{\prime}n}}<0$ can be found in Appendix B. We randomly choose a vector $\bm{h}\in\mathbb{R}^{M\times 1}$ with elements not all 00. For $\bm{\Lambda}_{n}$, it is easy to see that $\bm{h}^{T}\bm{\Lambda}_{n}\bm{h}=\sum_{m}\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}(h^{m})^{2}<0$. Furthermore, $\bm{h}^{T}\bm{H}_{n}\bm{h}=\frac{-\mu_{n}}{(1+\sum_{m}V_{mn})^{2}}\left(\sum_{m}h^{m}V_{mn}^{\prime}\right)^{2}<0$. Thus, we have $\bm{h}^{T}(\bm{\Lambda}_{n}+\bm{H_{n}})\bm{h}=\bm{h}^{T}\bm{\Lambda}_{n}\bm{h}+\bm{h}^{T}\bm{H}_{n}\bm{h}<0$, indicating that the utility function $\Psi_{n}$ is strictly concave. So far, the uniqueness of the Nash equilibrium [37] solution to the multi-MSP rewarding game $\Omega$ has been proved. ∎
1 Input: Action threshold $[p_{min},p_{max}]$ and number of episodes $\mathcal{T}$;
@@
3 # Trading process using MDDR:
@@ 15       Clear the episode buffer $\mathcal{D}_{n}$ for each MSP agent $n$;
17Return: Reward decisions $\bm{p}$ of MSPs and resource allocation $[\bm{f},\bm{B}]$ of MUs;
18 # FL training guided by trading results within time $T$:
@@
21             Based on the trading results, utilize $f_{mn}$ and $B_{mn}$ to perform the local FL task for MSP $n$, and then upload the model to MSP $n$.
@@
23      Aggregate received local models and issue an updated global model;
@@ 24       Supply the updated model to the AR engine;
Utilizing Theorem 1 and Theorem 2, we prove the existence and uniqueness of the hierarchical equilibrium in the proposed EPEC model. Concretely, each MU has a unique best-response resource allocation strategy in reaction to any given MSP rewards, while each MSP determines a unique optimal pricing to maximize its own utility given the rewards of other MSPs.
@@ VI Dynamic Reward Strategy for MSPs
According to the EPEC proven above, it is possible to obtain an optimal solution. However, the optimal approach faces the following practical challenges: (i) In a time-varying network, the optimal approach is time-consuming due to the complexity of (12), which is a non-linear problem with a complicated structure. Additionally, the reward decisions of MSPs are tightly coupled. (ii) MSPs must have prior knowledge of private information about MUs (e.g., cost factors) to determine their rewards, which raises privacy concerns for MUs. To address these challenges, we further formulate the multi-MSP rewarding game as a multi-agent Markov decision process (MAMDP) [35], which adapts to dynamic channel conditions. In this framework, each MSP is modeled as an individual agent that makes intelligent reward decisions.
@@ VI-A MAMDP for Multi-MSP Rewarding Game $\Omega$
We train the reward model based on the state-of-the-art policy gradient method proximal policy optimization (PPO) for the reasons described in [38]. The $MAMDP=\left\langle\mathcal{S}_{n},\mathcal{A}_{n},\mathcal{P}_{n},\mathcal{U}_{n}\right\rangle$ for multi-MSP rewarding game is composed of state space $\mathcal{S}_{n}\triangleq\left\{s_{n}\right\}$, action space $\mathcal{A}_{n}\triangleq\left\{\bm{p}_{n}\right\}$, state transition probability $\mathcal{P}_{n}\triangleq\left\{P_{n}\right\}$, and utility $\mathcal{U}_{n}\triangleq\left\{\Psi_{n}\right\}$. For the MSP agent, the local observation contains the channel information and the responses of MUs, all of which are captured in the IoM of (2). Thus, we set the local observation of MSP $n$ at the $k$-th stage game defined as $s_{n}(k)=[V_{1n}(k),V_{2n}(k),\ldots,V_{mn}(k)]$, and the state of the environment is $\mathcal{S}(k)=[s_{1}(k),\ldots,s_{n}(k)]$. At the $k$-th stage game, MSP $n$ observes a state $s_{n}(k)$ and determines an action $\bm{p}_{n}(k)$ within $[p_{min},p_{max}]$. When an action $\bm{p}_{n}(k)$ is applied to state $s_{n}(k)$, the agent $n$ receives a utility $\mathcal{U}_{n}(k)$ from the environment. Taking into account the competition among agents, we define the reward $\mathcal{U}_{n}$ for each MSP by (12). The state transition probability $P(s_{n}(k+1)\mid s_{n}(k),\bm{p}_{n}(k))$ leads to the new state $s_{n}(k+1)$ after executing an action $\bm{p}_{n}(k)$ at the state $s_{n}(k)$.
@@ VI-B The Multi-agent DRL-based Dynamic Reward (MDDR)
Trading process: At the start of the game, each MSP $n$ initializes its actor $\bm{\alpha}_{n}$ and critic network $\bm{\beta}_{n}$, and episode buffer $\mathcal{D}_{n}$ for each MSP agent $n$ (Line 2). Each agent $n$ feeds the observed information into the policy network $\bm{\alpha}_{n}$ to derive its reward policy (Lines 7-8). Subsequently, each MSP agent $n$ broadcasts its reward decisions, interacts with the environment, receives feedback from the MUs, calculates its utility, and stores the experience in the episode buffer $\mathcal{D}_{n}$ (Lines 9-12). The experience is collected through this interactive process (Lines 7-12) until the buffer $\mathcal{D}_{n}$ is full. Once the buffer is full, the critic network $\bm{\beta}_{n}$ evaluates the actor network $\bm{\alpha}_{n}$ based on the collected experience and updates it accordingly (Line 14). Finally, the episode buffer is cleared in preparation for the next episode (Line 15). The game will terminate when the maximum number of training episodes is reached and then the trading results will be obtained.
@@
VI-C The Complexity and Convergence of MDDR
The complexity of MDDR is mainly influenced by the structure of the actor-critic network. According to [39], the total computational complexity of the fully connected layer can be expressed as the number of multiplications $\mathcal{O}(\sum_{u=1}^{U}\varrho_{u}\cdot\varrho_{u-1})$, where $\varrho_{u}$ is the number of neurons in the $u$-th hidden layer. In this work, each agent has its actor-critic network, structured as $[5,256,10]$ and $[5,256,1]$, respectively. Therefore, its complexity can be expressed as $\mathcal{O}(5\times 256+256\times 10+5\times 256+256\times 1)$ without burden on the MSP. After several attempts, the MDDR can achieve convergence [40] (as per Fig. 7 and Table II) with a suitable set of hyper-parameters.
TABLE II: The Utilities of three reward approaches
@@ VII-A Simulation Settings
[…]
@@ VII-B Benchmark Schemes
  • x_based: Inspired by the incentive mechanism of the Stackelberg game for federated learning [31], the MU allocates computational and communication resources based on the amount of data involved in local training.
@@
TABLE III: Time comparison of five schemes for achieving the same performance.

### v3

@@ IV-C Multi-MSP Multi-MU Game as EPEC
As shown in Fig. 6, MUs and MSPs negotiate allocation strategies of computational and communication resources and reward decisions to maximize their benefits. In the upper level, MSPs determine the rewards they are willing to offer by considering their cost, the responses of MUs, and the decisions of other MSPs. In the lower level, each MU $m$ gives the optimal allocation response for computational and communication resources by considering its resource constraints, cost, and rewards from different MSPs.
The objective of EPEC is to find the equilibria at two levels, i.e., the point at which the MSPs’ (leaders’) utilities are maximized given that the MUs (followers) will choose their best responses. For the proposed EPEC, the equilibria at two levels are defined as follows.
@@
$\Psi_{n}\left(\bm{f}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{B}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{p}_{n}^{*},\bm{p}_{-n}^{*}\right)$
$\geq\Psi_{n}\left(\bm{f}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{B}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{p}_{n},\bm{p}_{-n}^{*}\right),\forall n\in\mathcal{N},$
where $\bm{p}_{-n}^{*}$ denotes the optimal reward vector for all MSPs except $n$.
@@
$\max_{p_{n}}\,\,\Psi_{n}=\psi_{n}-\sum_{m=1}^{M}\bigg[p_{mn}I_{mn}\bigg(\frac{1}{2}\tau_{n}-\frac{x_{m}\tau_{n}\log(1/\theta_{m})}{f_{mn}^{*}}$ (14)
$\qquad\qquad\qquad\qquad\qquad-\frac{b_{mn}}{B_{mn}^{*}\log_{2}(1+\varsigma_{mn})}\bigg)\bigg],$
$\mathrm{s.t.}\ \left\{\begin{aligned} &p_{mn}>0,\ n\in\mathcal{N},\ m\in\mathcal{M},\\ &(\bm{f}_{m}^{*},\bm{B}_{m}^{*})=\arg\max\Phi_{m}(\bm{f}_{m},\bm{B}_{m}),\\ &\qquad\qquad\quad\quad\mathrm{s.t.}\ C1,C2,C3.\end{aligned}\right$.
@@ V-A Lower Level: Optimal Resource Allocation for MUs
Since $\bm{H}^{f}_{m}$ and $\bm{H}^{B}_{m}$ are diagonal matrices with strictly negative diagonal elements, they are negative definite. Consequently, the block-diagonal Hessian $\bm{H}_{m}$ is negative definite, implying $\Phi_{m}(\bm{f}_{m},\bm{B}_{m})$ is strictly concave and continuous. Moreover, the objective function $\Phi_{m}$ tends to -$\infty$ as any variable approaches the boundary of the feasible set (e.g., $f_{mn}\to 0^{+}$ and $B_{mn}\to 0^{+}$, or $f_{mn}\to\infty$ and $B_{mn}\to\infty$). Hence, the maximum is necessarily attained in the interior. Combined with strict concavity over the convex feasible region defined by constraints C1-C3, this ensures that Problem 1 admits a unique global optimum $(\bm{f}_{m}^{*},\bm{B}_{m}^{*})$ (see Section IV in [38]). ∎
Consequently, MU $m$ has a unique best-response resource allocation for each MSP by solving Problem 1. From the Karush–Kuhn–Tucker (KKT) conditions, we derive the following proposition.
@@
$f_{mn}^{*}=\begin{cases}\sqrt{\frac{p_{mn}I_{mn}x_{m}\tau_{n}}{c_{m}^{f}}},&p_{mn}\geq\frac{F_{mn}^{2}}{I_{mn}},\\ F_{mn}\sqrt{\frac{x_{m}\tau_{n}}{c_{m}^{f}}},&0<p_{mn}<\frac{F_{mn}^{2}}{I_{mn}},\\ 0,&\text{otherwise},\end{cases}$ (18)
@@
$B_{mn}^{*}=\begin{cases}\sqrt{\frac{p_{mn}I_{mn}b_{mn}}{c_{m}^{B}\log_{2}(1+\varsigma_{mn})}},&p_{mn}\geq\frac{F_{mn}^{2}}{I_{mn}},\\ F_{mn}\sqrt{\frac{b_{mn}}{c_{m}^{B}\log_{2}(1+\varsigma_{mn})}},&0<p_{mn}<\frac{F_{mn}^{2}}{I_{mn}},\\ 0,&\text{otherwise},\end{cases}$ (19)
where $F_{mn}=2\log(1/\theta_{m})\sqrt{\frac{x_{m}c_{m}^{f}}{\tau_{n}}}+\frac{2\sqrt{b_{mn}c_{m}^{B}}}{\tau_{n}\sqrt{\log_{2}(1+\varsigma_{mn})}}$. For the detailed proof of Proposition 1, please refer to Appendix A.
@@ V-B Upper Level: Optimal Reward Equilibrium among MSPs
There exists a unique Nash equilibrium in the multi-MSP rewarding game $\Omega$, ensuring a unique optimal reward decision profile $\{\bm{p}_{n}^{*}\}_{n\in\mathcal{N}}$.
@@
We randomly choose a vector $\bm{h}\in\mathbb{R}^{M\times 1}$ with elements not all 0. Then, we have $\bm{h}^{T}(\bm{\Lambda}_{n}+\bm{H_{n}})\bm{h}=\sum_{m=1}^{M}(h^{m})^{2}(\frac{\mu_{n}V_{mn}^{\prime\prime}}{1+\sum_{m=1}^{M}V_{mn}}-2V_{mn}^{\prime}-p_{mn}V_{mn}^{\prime\prime})-\frac{\mu_{n}}{(1+\sum_{m=1}^{M}V_{mn})^{2}}(\sum_{m=1}^{M}h^{m}V_{mn}^{\prime})^{2}<0$, indicating that the utility function $\Psi_{n}$ is strictly concave. According to  [39], there exists a unique Nash equilibrium in the multi-MSP rewarding game $\Omega$. The proof of $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}<0$, $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}\partial p_{m^{\prime}n}}<0$, and $\bm{h}^{T}(\bm{\Lambda}_{n}+\bm{H_{n}})\bm{h}<0$ can be found in Appendix B. ∎
1Input: Action threshold $p_{min},p_{max}]$ and number of episodes $\mathcal{T}$;
@@
3
41ex# Trading process using MDDR:
@@ 16   Clear the episode buffer $\mathcal{D}_{n}$ for each MSP agent $n$;
18Return: Reward decisions $\bm{p}$ of MSPs and resource allocation $\bm{f},\bm{B}]$ of MUs;
19
201ex# FL training guided by trading results within time $T$:
@@
23      Based on the trading results, utilize $f_{mn}$ and $B_{mn}$ to perform the local FL task for MSP $n$, and then upload the local updates to MSP $n$.
@@
25  Aggregate received local updates and issue an updated global model;
@@ 26   Supply the updated model to the AR engine;
Utilizing Theorem 1 and Theorem 2, we prove the existence and uniqueness of the hierarchical equilibrium in the proposed EPEC model. Concretely, each MU has a unique best-response resource allocation strategy in reaction to any given MSP rewards. At the same time, each MSP determines unique optimal pricing to maximize its own utility, given the rewards of other MSPs.
@@ VI Dynamic Reward Strategy for MSPs
According to the EPEC proven above, it is possible to obtain an optimal solution. However, the optimal approach faces the following practical challenges: (i) In a time-varying network, the optimal approach is time-consuming due to the complexity of (12), which is a non-linear problem with a complicated structure. Additionally, the reward decisions of MSPs are tightly coupled. (ii) MSPs must have prior knowledge of private information about MUs (e.g., cost factors) to determine their rewards, which raises privacy concerns for MUs. To address these challenges, we further formulate the multi-MSP rewarding game as a multi-agent Markov decision process (MAMDP) [36], adapting to dynamic channel conditions. In this framework, each MSP is modeled as an individual agent that makes intelligent reward decisions.
@@ VI-A MAMDP for Multi-MSP Rewarding Game $\Omega$
We train the reward model based on the state-of-the-art policy gradient method proximal policy optimization (PPO) for the reasons described in [40]. The $MAMDP=\left\langle\mathcal{S}_{n},\mathcal{A}_{n},\mathcal{P}_{n},\mathcal{U}_{n}\right\rangle$ for multi-MSP rewarding game is composed of state space $\mathcal{S}_{n}\triangleq\left\{s_{n}\right\}$, action space $\mathcal{A}_{n}\triangleq\left\{\bm{p}_{n}\right\}$, state transition probability $\mathcal{P}_{n}\triangleq\left\{P_{n}\right\}$, and utility $\mathcal{U}_{n}\triangleq\left\{\Psi_{n}\right\}$. For the MSP agent, the local observation contains the channel information and the responses of MUs, all of which are captured in the IoM of (2). Thus, we set the local observation of MSP $n$ at the $k$-th stage game defined as $s_{n}(k)=[V_{1n}(k),V_{2n}(k),\ldots,V_{mn}(k)]$, and the state of the environment is $\mathcal{S}(k)=[s_{1}(k),\ldots,s_{n}(k)]$. At the $k$-th stage game, MSP $n$ observes a state $s_{n}(k)$ and determines an action $\bm{p}_{n}(k)$ within $p_{min},p_{max}]$. When an action $\bm{p}_{n}(k)$ is applied to state $s_{n}(k)$, the agent $n$ receives a utility $\mathcal{U}_{n}(k)$ from the environment. Taking into account the competition among agents, we define the reward $\mathcal{U}_{n}$ for each MSP by (12). The state transition probability $P(s_{n}(k+1)\mid s_{n}(k),\bm{p}_{n}(k))$ leads to the new state $s_{n}(k+1)$ after executing an action $\bm{p}_{n}(k)$ at the state $s_{n}(k)$.
@@ VI-B The Multi-agent DRL-based Dynamic Reward (MDDR)
Trading process: At the start of the game, each MSP $n$ initializes its actor $\bm{\alpha}_{n}$ and critic network $\bm{\beta}_{n}$, and episode buffer $\mathcal{D}_{n}$ for each MSP agent $n$ (Line 2). Each agent $n$ feeds the observed information into the policy network $\bm{\alpha}_{n}$ to derive its reward policy (Lines 7-8). Subsequently, each MSP agent $n$ broadcasts its reward decisions, interacts with the environment, receives feedback from the MUs, calculates its utility, and stores the experience in the episode buffer $\mathcal{D}_{n}$ (Lines 9-12). The experience is collected through this interactive process (Lines 7-12) until the buffer $\mathcal{D}_{n}$ is full. Once the buffer is full, the critic network $\bm{\beta}_{n}$ evaluates the actor network $\bm{\alpha}_{n}$ based on the collected experience and updates it accordingly (Line 14). Finally, the episode buffer is cleared in preparation for the next episode (Line 15). The game terminates once the maximum number of training episodes is reached, after which the trading results are finalized.
@@
TABLE III: The Utilities of three reward approaches
@@ VII-A Simulation Settings
We conduct experiments on the MNIST dataset with ResNet-18 and the GTSRB dataset with Faster R-CNN. The GTSRB was recorded during daytime driving on different types of roads in Germany for vehicle AR scenario testing [41]. We set the number of MUs as $M=5$ and the number of MSPs as $N=3$. The configurations of the maximum computational and communication resource for each MU are randomly generated from the range [3,5][3,5]GHz and [1,4][1,4]MHz [42]. For local training of MUs, we adopt the SGD optimizer with $\text{momentum}=0.9$ and $\text{learning rate}=0.001$. For model aggregation, MSPs adopt the Federated Averaging (FedAvg) algorithm, applying equal weights to all participants’ updates. The trading guidance time $T$ for MNIST and GTSRB are $30$s and $1,200$s, respectively. All baseline methods are configured with the same settings for fair comparison.
@@ VII-B Benchmark Schemes
  • x_based: Inspired by the incentive mechanism of the Stackelberg game for federated learning [32], the MU allocates its computational and communication resources based on the amount of data involved in local training.
@@
VII-C Performance Evaluation of MDDR
First, we consider a relatively static communication environment (fixed channel conditions) to show the effectiveness of MDDR compared to the optimal solution and the general multi-agent depth approximation policy optimization (MAPPO) approach. We trained our proposed MDDR and MAPPO until convergence, and the results are shown in Fig. 7 and Table III. We find that both MDDR and MAPPO achieve good results, but MAPPO slightly outperforms MDDR. This is attributed to MAPPO’s centrally controlled yet distributed execution architecture [44] that can utilize more information for decision making. However, this approach is not suitable for non-cooperation due to the need to share parameters among MSPs.
For the fully distributed MDDR, it is more suitable for non-cooperative scenarios. The reason is that in MDDR, each agent has a self-contained actor-critic network, and information is not required to be shared. This suggests that the approach can achieve considerable utility while preserving privacy and saving communication costs significantly. Then, we deploy the MDDR in dynamic communication environments where channel conditions change over time. The trading behaviors of each MU and MSP are illustrated in Figs. 8(a), 8(b), 8(c) and 8(d), reflecting that the MDDR can converge rapidly and show good stability in dynamic network environments.
VII-D Benchmark Comparison
TABLE IV: Time comparison of five schemes for achieving the same performance